# SF-R You Sure? The Conflicting Role of Star Formation Rates in Constraining the Evolution of Milky Way Analogues in Cosmological Simulations

### v1

@@ A.1 Chemical enrichment
Finally, we note that the observed abundance plane for the MW as shown in Figure 12 was afterwards fed into the existing UMAP embedding and is indicated in the MW-ness UMAP projections in the right column of Figure 13 as a black star. The true abundance plane does not fall within the MWA region for either simulation. In EAGLE, the MW falls on the opposite end of the projection space as the MWAs, and is shown to be in the head of the tadpole. TNG does a bit better, with the MW appearing at the very top of the head of the duck, towards the edge of the MWA region. Interestingly, the position of the MW in the projection space corresponds to much too low SM and SFR in the EAGLE simulation, but slightly too high SM and SFR in the TNG simulation. It seems that neither simulation is accurately reproducing the MW’s abundance plane, despite both being self-consistent in producing MWA abundance planes. The reasoning behind this discrepancy between simulation and observation is beyond the scope of this work, and should be addressed in future research.
@@ A.2 B/T in EAGLE
Despite the skepticism brought on by the small sample size, we present the SFH and SMH of the MWAs selected on SM, SFR, and B/T (magenta) compared to those selected only on SM and SFR (green) in Figure 15. Generally, the SM-SFR-B/T MWAs peak higher in SFR and assemble their SM quicker than the original MWAs, but the overall profiles of both the SFH and SMH are consistent with the SM-SFR MWAs, if only on the upper end. We caution again that no strong conclusions can be drawn with such a small sample size, but note that the inclusion of B/T as a selection parameter did not drastically change our results. Rather, it only served to introduce additional complications with no added benefit, and so, B/T may not be a defining characteristic of MWAs – at least not in EAGLE.

### v2

@@ A.1 Chemical enrichment
Finally, we note that the observed abundance plane for the MW as shown in Figure 12 was afterwards fed into the existing UMAP embedding and is indicated in the MW-ness UMAP projections in the right column of Figure 13 as a black star. The true abundance plane does not fall within the MWA region for either simulation. In EAGLE, the MW falls on the opposite end of the projection space as the MWAs, and is shown to be in the head of the tadpole. TNG does a bit better, with the MW appearing at the very top of the head of the duck, towards the edge of the MWA region. Interestingly, the position of the MW in the projection space corresponds to much too low SM and SFR in the EAGLE simulation, but slightly too high SM and SFR in the TNG simulation. It seems that neither simulation is accurately reproducing the MW’s abundance plane, despite both being self-consistent in producing MWA -abundance planes. The reasoning behind this discrepancy between simulation and observation is beyond the scope of this work, and should be addressed in future research.
@@ A.2 B/T in EAGLE
Despite the skepticism brought on by the small sample size, we present the SFH and SMH of the MWAs selected on SM, SFR, and B/T (purple) compared to those selected only on SM and SFR (green) in Figure 15. Generally, the SM-SFR-B/T MWAs peak higher in SFR and assemble their SM quicker than the original MWAs, but the overall profiles of both the SFH and SMH are consistent with the SM-SFR MWAs, if only on the upper end. The fact that the inclusion of B/T seems to preferentially select earlier assembly histories is in agreement with recent studies finding early disk formation of the MW (Belokurov & Kravtsov, 2022; Rix et al., 2022; Conroy et al., 2022; Dillamore et al., 2022; Semenov et al., 2024a). We caution again that no strong conclusions can be drawn with such a small sample size, but note that the inclusion of B/T as a selection parameter did not drastically change our results. Rather, it only served to introduce additional complications with no added benefit, and so, B/T may not be a defining characteristic of MWAs – at least not in EAGLE.